A wire of length 7 π‘š and radius 2.37 π‘šπ‘š is stretched by a load of 12370 π‘”π‘š and

increased by 1.78% of the original length. Find the stress, strain, and Young’s

modulus of the material of the wire.



1
Expert's answer
2021-12-06T09:48:48-0500
"\sigma=\frac{F}{A}=\frac {mg}{\pi r^2}=6.87\text{ MPa}.\\\space\\
\epsilon=1.78\%/100\%=0.0178.\\\space\\
E=\frac{\sigma}{\epsilon}=386\text{ MPa}."